# Scientific Literature Hierarchy Construction and Organization

### v3

@@ 5.3 Experiment Design
We conduct few sets of experiments to evaluate our method (Scychic, §4.1) against the baselines (§4.2, 4.3) following the proposed evaluation protocol. Hyperparameter details for Scychic are provided in §G. The experiments are structured as follows: (1) We begin by comparing all methods on “topic” contributions, the simplest contribution type (Table 5.3). Due to the high computational cost, LLM-based baselines are evaluated only in this setting. (2) We then evaluate performance on more complex contributions (problem, solution, and results) using both SciPile and SciPileLarge to test scalability (Table 4). The results tables also report LLM Cost (average number of input tokens and LLM calls) and Hierarchy Structure (depth and branching factor per node).
@@ 5.4 Empirical Findings
LLM-based baselines can be far more expensive than Scychic. While fLMSci slightly outperforms Scychic in accuracy, it does so at the cost of a massive increase in LLM calls—making it impractical for large-scale use. As a result, despite its strong performance, fLMSci (incremental) simply doesn’t scale.
Scychic scales to larger paper corpus. For our scalability experiments, we evaluated Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type. Due to the significant increase ($\times 5$) in corpus size, we implemented a 4-layer hierarchy instead of the 3-layer structure used previously. Notably, Scychic achieved even higher L1-Acc (86.5%) on SciPileLarge compared to our smaller dataset SciPile. This improvement likely stems from the enhanced quality of our expanded dataset, which has more strict filtering mechanisms. While the Strict-Acc showed a minor decrease compared to results on SciPile, it remained at a satisfactory level. Collectively, these results provide compelling evidence that our method scales successfully to substantially larger paper corpora.
@@ 5.6 Sample Visualization of the Hierarchy
The reader might be curious to see the resulting hierarchies. In §I we show a slice of the final hierarchy generated by Scychic on the SciPileLarge  dataset. The original hierarchy has 4 levels, use papers’ problem contribution. Due to space constraints, this slice shows only two levels of clusters above the individual papers.
@@ Conclusions:
Acknowledgments
This work is supported by ONR grant (N0001424-1-2089) and Defense Advance Research Projects Agency (DARPA) under Contract No. HR001125C0304. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of DARPA. We sincerely thank Jiefu Ou and the broader JHU CLSP community for discussions and inspiration.
@@ Appendix B Evaluation Framework
One question is, which LLM should we use for evaluation? We chose Qwen2.5-32b-instruct for its strong instruction-following capabilities. In pilot experiments, Qwen showed a high consistency against GPT4 score, compared to other open-weight models. Here’s a summary of that experiment: We evaluated one of the hierarchies produced by Scychic using different models, including GPT-4. Assuming GPT-4 has the highest accuracy, we sought alternative models with the greatest consistency against it, as frequent evaluations with GPT4 are costly. Fig.5 presents the results. As it can be observed, Llama has the highest agreement, but we suspect bias since the hierarchy was also constructed with Llama. To avoid this, we selected the next best model, Qwen2.5-32b-instruct, for evaluation.
@@ Appendix C Extracting Paper Contributions
Below are prompts and examples for extracting different contributions (problem, solution, result and topics) from papers’ titles and abstracts. we utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions.
@@ Appendix D Compiling a Seed Hierarchy
This section includes the pipeline and prompts used for fLMSci (parallel) and fLMSci(incremental).
@@ Appendix G Hyperparameters of Scychic
Here shows the models and hyperparameters we use for the experiments mentioned in §5.3. We utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions. For summarizer, we use We choose Llama-3.3-70B-InstructGrattafiori et al. (2024) for its superiority of following instructions among open-source models, and use gte-Qwen2-7B-instruct as our embedder. For clustering algorithm, we apply k-means clustering. The number of clusters for each layer is (6, 40, 276) when conducting experiments on SciPile ($2K$ papers), and (6, 40, 276, 1250) when on SciPileLarge ($10k$ papers).
@@ H.3 Experiments of Prompt Engineering
Appendix I Demonstration of Hierarchy
Below is a snippet of our final hierarchy result.

### v4

@@ 5.3 Experiment Design
We conduct a series of experiments to evaluate our method (Scychic, §4.1) against the baselines (§4.2, 4.3) using the proposed evaluation protocol. Hyperparameter settings for Scychic are detailed in §G. The experiments are organized as follows: (1) We first compare all methods on the simplest contribution type (“topic”) in Table 5.3. Due to the high computational cost, LLM-based baselines are evaluated only in this setting. (2) We then evaluate performance on more complex contributions (problem, solution, and results) using both SciPile and SciPileLarge to test scalability (Table 4). Each results table also reports LLM Cost (average input tokens and number of calls) and Hierarchy Structure (average depth and branching factor).
@@ 5.4 Empirical Findings
LLM-based approaches can be expensive. While fLMSci slightly outperforms Scychic in accuracy, it does so at the cost of a massive increase in LLM calls—making it impractical for large-scale use. As a result, despite its strong performance, fLMSci (incremental) simply doesn’t scale.
Scychic scales to larger paper corpus. For our scalability experiments, we evaluate Scychic on our larger $10K$ paper dataset SciPileLarge, using the problem statement contribution type. Due to the significant increase ($\times 5$) in corpus size, we extend the hierarchy to four layers (versus three previously). Notably, Scychic achieved even higher L1-Acc (86.5%) on SciPileLarge compared to our smaller dataset SciPile. This improvement likely stems from the enhanced quality of our expanded dataset, which has more strict filtering mechanisms. While the Strict-Acc showed a minor decrease compared to results on SciPile, it remained at a satisfactory level. Collectively, these results provide compelling evidence that our method scales successfully to substantially larger paper corpora.
@@ 5.5 Additional Analyses
Quality diagnostics confirm the reliability of the hierarchies. We further analyze cluster coherence by examining citation patterns within and across clusters. Out of 3,056 total citations, 2,587 (84.7%) occur between papers in the same cluster, while the remaining 469 (15.3%) are inter-cluster citations. The visualization and examples of inter-cluster citations can be found in §I.
@@ 5.6 Sample Visualization of the Hierarchy
The reader might be curious to see the resulting hierarchies. In §J we show a slice of the final hierarchy generated by Scychic on the SciPileLarge  dataset. The original hierarchy has 4 levels, use papers’ problem contribution. Due to space constraints, this slice shows only two levels of clusters above the individual papers.
@@ Appendix A Additional Related Work
Comparison with existing hierarchical systems:
Web of Science maintains a flat (one-level) collection of  250 research fields, which is useful for categorization. Given its flat structure, it is not a hierarchical structure. There are no parent-child relationships or summaries connecting broader and narrower concepts. These are best understood as only labels, not nodes in a multi-level taxonomy. Scopus uses a fixed-depth (2-layer) hierarchy based on research field names (ASJC Codes). Importantly, these codes are assigned at the journal level rather than to individual papers. Papers inherit classifications from their publishing journals, meaning the hierarchy is not derived from the actual paper content. PubMed MeSH terms provides hierarchical labeling for PubMed publications, but it functions at the level of keywords (few tokens) rather than leveraging the full richness of natural language from science papers. Crucially, it is organized around a fixed set of controlled terms rather than the actual semantic content of the papers, limiting its suitability for constructing dynamic or corpus-specific hierarchies. Additionally, because MeSH is manually curated, it introduces indexing delays—papers are only labeled after publication—and is subject to human bias, as noted by  Hadfield (2020). Microsoft Academic Graph (MAG), though discontinued in 2021, offered a rich graph-based structure connecting papers and authors. Its hierarchical classification derived primarily from citation patterns and machine learning clustering rather than semantic paper content, which limited cluster interpretability.
@@ Appendix B Evaluation Framework
B.1 Pilot Experiment for Evaluator Choice
One question is, which LLM should we use for evaluation? As discussed in §5.2, we chose Qwen2.5-32b-instruct for its strong instruction-following capabilities. In pilot experiments, Qwen showed a high consistency against GPT4 score, compared to other open-weight models. Here’s a summary of that experiment: We evaluated one of the hierarchies produced by Scychic using different models, including GPT-4. Assuming GPT-4 has the highest accuracy, we sought alternative models with the greatest consistency against it, as frequent evaluations with GPT4 are costly. Fig.5 presents the results. As it can be observed, Llama has the highest agreement, but we suspect bias since the hierarchy was also constructed with Llama. To avoid this, we selected the next best model, Qwen2.5-32b-instruct, for evaluation.
@@ B.1 Pilot Experiment for Evaluator Choice
B.2 Validation of LLM-based Evaluation
As we discuss in §5.2, to validate our evaluation framework, a Computer Science PhD student analyzes 200 error cases (50 cases per layer). For each case, the annotator determines whether the error comes from the LLM evaluator or from the hierarchy itself. The analysis reveals three types of cases. First, only 9 cases (4.5%) are clear evaluator errors. Second, in 39 cases (18.5%), both the evaluator’s choice and the hierarchy path are reasonable, which is expected for interdisciplinary works. Third, in the remaining 152 cases (77%), the evaluator agrees with the human annotator. These results confirm the reliability of our LLM-based evaluation approach.
To further validate our LLM-based evaluation approach, we downloaded the annotations from the Open Research Knowledge Graph (ORKG). On this website, papers are curated entirely by volunteers who are strongly familiar with the topics of the papers. We use the subset of the ORKG data focused on the Engineering domain. This led to a collection of 4.4K papers that are organized in a 2-layer hierarchy. Treating this data as a high-quality hierarchy, the question is whether our evaluation would assign it a high score. We ran our evaluation experiment with Qwen2.5-32B-Instruct as the LLM-as-a-judge. Similar to our setup from the paper, we use paper title/abstract as queries, and require the evaluator to traverse the hierarchy by incrementally making the most appropriate choice between all possible cluster candidates. Our results show that the evaluator model has an accuracy of 83% (i.e., in 83% of the runs it identified the correct paper). This indicates that our evaluation metric is able to assign a high score to a good hierarchy.
@@ Appendix C Extracting Paper Contributions
As we discuss in §3.2, below are prompts and examples for extracting different contributions (problem, solution, result and topics) from papers’ titles and abstracts. we utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions.
@@ Appendix D Compiling a Seed Hierarchy
This section includes the pipeline and prompts used for fLMSci (parallel) and fLMSci(incremental) from §4.3.
@@ Appendix G Hyperparameters of Scychic
Here shows the models and hyperparameters we use for the experiments mentioned in §5.3. We utilize the GPT-4o model (gpt-4o-2024-08-06) to generate all contribution extractions along with detailed rationales explaining the extraction decisions. For summarizer, we use Llama-3.3-70B-InstructGrattafiori et al. (2024) for its superiority of following instructions among open-source models, and use gte-Qwen2-7B-instruct as our embedder. For clustering algorithm, we apply k-means clustering. The number of clusters for each layer is (6, 40, 276) when conducting experiments on SciPile ($2K$ papers), and (6, 40, 276, 1250) when on SciPileLarge ($10k$ papers).
@@ H.3 Experiments of Prompt Engineering
Appendix I Visualization and Examples of Inter-Cluster Citations
Below is the figure of comparing inter- and intra-cluster citation counts mentioned in §$5.$
Figure 13: Comparison of inter(green)- and intra(blue)-cluster citation counts. 84.7% citations are between papers in the same top layer cluster, and the rest inter-cluster citations are mostly theory-to-application works, which proves the reliability of Scychic.
Below are examples of citations between papers in different top-layer clusters. These examples show that many inter-cluster citations represent theory-to-application connections, while the last row illustrates cross-disciplinary citations between research fields. Importantly, all papers involved are correctly categorized—the inter-cluster citations reflect legitimate relationships rather than classification errors.
Citing Paper Cited Paper
Rationale: AI research grounding in foundational cognitive science theory
[1pt/2pt]
[…]
Rationale: Theory-to-application for THz photonics
[1pt/2pt]
[…]
Rationale: Hardware implementation citing quantum network theory
[1pt/2pt]
[…]
Rationale: Manufacturing citing characterization techniques
[1pt/2pt]
[…]
Rationale: Cross-disciplinary bridge between biology and quantum physics
[1pt/2pt]
[…]
Table 11: Examples of cross-cluster citations. Each row shows the citing paper, the cited paper, their cluster names, and the citation rationale.
Appendix J Demonstration of Hierarchy
Below is a snippet of our final hierarchy result as mentioned in §5.6.